OpenStudy (anonymous):

Can anyone help me on this one? Need help on figuring this answer: ln(3) - 2ln(4+8) I got 3.9 and it was wrong...please show your work, so I can understand what I did wrong

OpenStudy (anonymous):

would it be negative 3.9?

OpenStudy (anonymous):

i did that too and it was wrong...i'm at a loss

OpenStudy (lalaly):

as i told u, i guess ln3 -ln12^2 =ln3/44 =-2.68

OpenStudy (anonymous):

sorry did not see your response...all that it said was 'I got'